# Efficient Tactile Perception with Soft Electrical Impedance Tomography and Pre-trained Transformer

### v1

@@ III-A3 Implementation
We designed the voltage encoder $\mathcal{E}_{vol}$ and decoder $\mathcal{D}_{vol}$ with customized scales and hyperparameters to match the characteristics of the E2IM input. To effectively capture information, the embedding dimension was optimized and set to 256, aligning with the resolution of the E2IM in $64\times 64$. The architecture adopts an asymmetric design to enhance the encoder’s feature extraction capacity while simplifying the decoder. Specifically, the voltage encoder $\mathcal{E}_{vol}$ consists of 12 Transformer layers, whereas the voltage decoder $\mathcal{D}_{vol}$ is deliberately kept shallower with only 2 Transformer layers. All Transformer layers are configured with 4 attention heads to maintain architectural consistency.
@@ III-A4 E2IM’s enhancement on pretraining
The study of MAE [37] has shown that mask sampling strategies significantly impact model performance, although the underlying reasons were not explored in depth. In MAE, an ablation study was conducted on highly semantic images from ImageNet[41], where the less semantic structural information masked, the easier the image could be reconstructed.
However, as depicted in Fig. 1, EIM represents a low-semantic representation of the voltage. In this context, applying large patch masking obscures key structural information, leading to poor reconstruction. Conversely, applying small patch masking reduces the representation quality, which hampers effective feature extraction. For example, $1\times 1$ masking will allow reconstruction through interpolation but limits the model’s ability to learn meaningful representations. Therefore, the minimum and optimal patch size $P$ for a $16\times 16$ EIM is $2$.
As we introduced before, to address this issue, we extended EIM from $16\times 16$ to $64\times 64$ and set the patch resolution to $4\times 4$, enhancing the model’s ability to learn and represent features effectively. To validate the effectiveness of the E2IM, we conducted ablation studies comparing its performance against the original EIM in terms of reconstructions in pertaining. As summarized in Table. I, MSE and Structural Similarity Index Measure (SSIM) were calculated using the mean values over 25,000 samples. The results demonstrate that E2IM achieves a 99.95% improvement in MSE and a 66.07% increase in SSIM over EIM, underscoring its superior representation quality.
@@ III-B1 Framework
$\arg\min_{\mathcal{E}_{tact},\mathcal{D}_{tact}}\mathcal{L}_{tact}\|\hat{\textbf{t}}-\textbf{t}\|^{2},$ (11)
@@ III-B2 Implementation
The tactile autoencoding model $\mathcal{D}_{tact}(\mathcal{E}_{tact}(\textbf{t}))$ is stacked with convolutional ResNet blocks. Each convolutional ResNet block consists of two ResNet layers followed by a convolution layer. These blocks perform downsampling or upsampling, with the number of channels in each layer ranging from [32, 64, 128].
@@
III-C Fine-tuning with limited volatge-tactile pairs
One of the primary challenges in learning-based tactile sensing lies in the scarcity of annotated real-world data for training. To address this limitation, we employ fine-tuning by integrating the pretrained voltage encoder $\mathcal{E}_{vol}$ and the pretrained tactile decoder $\mathcal{D}_{tact}$ into a unified model. This model is then refined using a limited dataset of voltage-tactile paired samples, enabling effective adaptation to the target task despite the constrained data availability. Due to the vast differences in scale between the small fine-tuning datasets (i.e., down to 2500 paired training sets) and large pretraining datasets (i.e., 450,000 unlabeled training sets), it often leads to over-fitting. To solve this problem, we propose strategies for few-shot fine-tuning for tiny datasets as follows:
@@ III-C1 Latent space compression
As outlined in Sections III-A and III-B, the deep neural network architecture of the autoencoder achieves a high compression ratio by encoding voltage and tactile data into a low-dimensional latent space. This compressed representation facilitates efficient data transformation and adaptation during the fine-tuning process.
@@ III-C2 Partial tuning
we observed significant differences in the number of parameters between the voltage encoder $9.54M$ and the tactile decoder $1.23M$. Therefore, as shown in Fig. 2, we freeze all parameters in the voltage encoder $\mathcal{E}_{vol}$ and only update the parameters in the tactile decoder $\mathcal{D}_{tact}$ during the fine-tuning stage, as shown in Fig.2. The latent variables from voltage measurement conditioning $\textbf{z}_{v}$ and tactile $\textbf{z}_{t}$ are connected via direct linear projection without any hidden layers:
@@
$\arg\min_{\mathcal{D}_{tact}}\mathcal{L}_{tact}\|\mathcal{D}_{tact}(\textbf{z}_{t})-\textbf{t}\|^{2},$ (16)
@@ IV-A Dataset
As shown in Fig. 3, the sensing region, measuring $100\times 100\,mm^{2}$, is divided into $48\times 48$ pixels with a background conductivity of 0.00312 S/m. Circular areas with diameters ranging from 7.5 mm to 27.5 mm are used to represent tactile interactions, emphasizing both position and intensity. The conductivity within these touch regions is randomly varied between 0.05 and 2 times the background value to simulate diverse touch effects. The dataset generated comprises 500,000 samples, with each of the five subsets containing 100,000 samples. Each sample includes one E2IM and corresponding Ground Truth (GT). These subsets correspond to different numbers of touch regions, ranging from 1 to 5 touch areas per sample. Each subset is further partitioned into training, validation, and testing datasets using an 18:1:1 ratio. This results in a total of 450,000 samples for training, 25,000 for validation, and 25,000 for testing. All data were generated using COMSOL Multiphysics and Matlab.
@@ IV-C Quantitative evaluation
Quantitative Metrics RE PSNR CC MSE (×$\times 10^{-2}$)
@@ V-A Sensor Fabrication
We designed a dual-conductivity layer tactile sensor combining a hydrogel layer and a carbon-black/graphite (CBG) conductive layer, leveraging the complementary properties of each material. Hydrogels are widely utilized in electronic skin applications due to their preferable electrical properties, flexibility and biocompatibility [44, 28]. The CBG layer with lower resistance improves EIT-based sensor performance.
The hydrogel precursor synthesis (Fig. 6a) was adapted from a formulation in [44], and optimized for tactile sensing. The hydrogel is primarily composed of acrylamide, a hydrophilic monomer known for polyacrylamide-based hydrogels, which retains water and ensures sensor functionality in humid conditions. The inclusion of lithium chloride, a common ionic conductor, further enhances the hydrogel’s ionic conductivity. The synthesis procedure is as follows:
  1. 1.
    The solution contained 21.5 weight percent (wt%) lithium chloride (Sigma-Aldrich), 8 wtacrylamide (Sigma-Aldrich), 32.27 wtdeionized water and 37 wtethylene glycol (Sigma-Aldrich) was prepared. Ethylene glycol maintains hydrogel pliability across a wide temperature range, essential for stable sensor performance.
  2. 2.
    The solution was stirred at 60°C and 700 rpm for two hours to achieve uniform mixing and dissolution.
  3. 3.
    To initiate a polyacrylamide-based network, 1 wtpolyacrylamide (Sigma-Aldrich) was introduced, providing mechanical stability and enhancing sensor durability.
  4. 4.
    The solution was stirred overnight at 60 ${}^{\circ}\mathrm{C}$ and 700 rpm to ensure complete dissolution and formation of a uniform polymer network.
  5. 5.
    A crosslinking agent, 0.15 wtN,N’-methylenebisacrylamide (Sigma-Aldrich), was added to control the mechanical strength of the gel. Additionally, 0.08 wt2-hydroxy-2-methylpropiophenone (Sigma-Aldrich), a photoinitiator, was added to enable UV curing of the hydrogel.
  6. 6.
    Finally, the mixture was shielded from light and stirred for an additional two hours at 60 ${}^{\circ}\mathrm{C}$ and 700 rpm to ensure complete homogenization.
The CBG conductive layer (Fig. 6c) was fabricated by combining carbon black and graphite. This blend leverages the high surface area and conductivity of carbon black along with the structural integrity and enhanced conductivity of graphite. The fabrication procedure is as follows:
  1. 1.
    3 g of carbon black (Imerys) and 2 g of graphite (Imerys) were dispersed in 23 g of Isopropyl Alcohol (IPA, RS Components) to create a uniform CBG mixture. IPA acts as a dispersing agent, ensuring even distribution of carbon black and graphite particles.
  2. 2.
    Ecoflex 00-30, a silicone elastomer chosen for its superior flexibility, was used as the substrate for the conductive layer. A 1:1 mixture of parts A and B of Ecoflex 00-30 (20 g each, Smooth-On Inc.) was stirred for 3 minutes to initiate cross-linking.
  3. 3.
    20 g of the silicone solution was added to the CBG mixture and stirred for three minutes, followed by degassing to remove any trapped air bubbles.
  4. 4.
    The remaining silicone solution was poured onto a glass plate, and a micrometer-adjustable applicator (VF1823, TQC Sheen) was used to achieve a uniform layer thickness of 0.1 mm. The film was allowed to degas for 10 minutes at room temperature before being cured in an oven at 100 ${}^{\circ}\mathrm{C}$ for 3 minutes.
  5. 5.
    The prepared CBG solution was applied over the cured silicone layer, again using the micrometer-adjustable applicator to maintain a uniform thickness of 0.1 mm. The film was degassed at room temperature for 10 minutes and subsequently cured in an oven at 100 ${}^{\circ}\mathrm{C}$ for another 3 minutes.
  6. 6.
    The cured composite film was cut into a 120 mm $\times$ 120 mm square using a laser cutter.
Finally, the hydrogel and CBG layers were integrated to create the tactile sensor (Fig. 6b). To ensure the sensor’s durability, the overall structure was encapsulated with a silicone layer to prevent moisture loss. The sensor dimensions are 120 $\times$ 120 $\times$ 10 mm3. The assembly procedure is as follows:
  1. 1.
    A 3D-printed mold, containing designated regions for the sensing area, cable channels, and electrode holes, was prepared.
  2. 2.
    Cables and electrodes were installed within the mold.
  3. 3.
    The back of the mold was sealed with a layer of silicone to minimize water loss from the hydrogel due to air exposure.
  4. 4.
    The hydrogel precursor solution was poured into the sensing area and subsequently cured under UV light (365 nm) for two hours to form a stable hydrogel layer.
  5. 5.
    The cured hydrogel layer was then combined with the CBG conductive layer, forming the primary sensing structure.
  6. 6.
    A layer of silicone was applied over the entire sensor assembly to provide environmental sealing and prevent external interference.
@@
The dual-layer design combines the flexibility and biocompatibility of the hydrogel with the enhanced conductivity and structural integrity of the CBG layer, maximizing signal response under applied pressure and making it well-suited for applications in electronic skin and human-machine interfaces.
@@ V-B Sensor Characterization
To evaluate the sensitivity of the developed sensor, we conducted an experiment where the sensor was subject to a continuously increasing localized pressure with strains ranging from 0 mm to 5 mm, as illustrated in Fig. 7. The results demonstrate a clear increase in the conductivity distribution as the compressive strain increases. Conductivity distribution images were generated for compressive strains of 1 mm, 2 mm, and 3 mm. At a compressive strain of 1 mm, the changes in conductivity are subtle yet detectable, while at 2 mm, the effect becomes more pronounced. These results demonstrate the sensor’s preferred sensitivity and capability for accurate and responsive tactile detection.
@@
To evaluate the sensor’s durability and repeatability, we performed a cycling test involving 300 pressing cycles with compressive strains from 2 mm to 5 mm. The results show that the sensor maintains stable output throughout the cycles, with consistent trends in signal changes (Fig. 8). During this test, a dynamic baseline approach was adopted to mitigate slight signal drifts attributed to the lagging effect of the sensor after prolonged pressing. These results confirm the sensor’s robustness and suitability for applications requiring long-term, repeated tactile interactions.
Figure 8: Sensor cycling test for 300 pressing cycles.
@@ VI-A Real-world Dataset
A diverse real-world dataset was created using 84 custom-fabricated touch tips produced by a 3D printer. The design of the touch points, including their positions and sizes, was randomly generated, with pressing depths precisely controlled between 2 mm and 5 mm. Data collection was performed using a three-axis linear robot (IGUS, DLE-RG-0001) to ensure precise control over the position and depth of the presses. An in-house developed EIT system was utilized for data acquisition as described in [46].
We collected a total of 5700 annotated samples, distributed as follows: 613 samples with 1 touch point, 1,063 with 2 touch points, 1,043 with 3 touch points, 1,004 with 4 touch points, 995 with 5 touch points, 493 with annular touch and 489 with L-shaped touch. Each sample consists of a 104-dimensional vector of voltage measurements, which is transformed into a 64x64 E2IM matrix and a 48x48 image representing the press map. The samples were randomly shuffled and split into three datasets in a 7:2:1 ratio for training, validation, and testing. To expedite training, we reused the pre-trained models from the simulation phase for both the PTET algorithm and the baseline DNN algorithm. This approach is justified because the pretraining model only captures the fundamental features of voltage measurements and pressure maps.
@@ VI-B Quantification of Reconstruction
Quantitative Metrics RE PSNR CC MSE (×$\times 10^{-2}$)
@@ VI-C Performance on Complex Touches
To evaluate the model’s effectiveness in tactile reconstruction during practical operations, we tested it with a series of complex touch scenarios. These included configurations involving multiple fingers, annular touch, annular touch with finger, and L-shaped touches. Fig. 10 shows the model’s ability to generalize effectively to diverse real-world touch interactions even without specific labels. These results demonstrate the model’s adaptability and robustness for practical applications. Supplementary video 1 shows the real-time tactile reconstruction performance on selected phantoms.
@@ VI-D Comparison of Different Training Datasets
CC 0.8633 0.4349 0.4455
MSE (×$\times 10^{-2}$) 0.1551 0.9069 0.8829
Best results are highlighted in bold.
@@
VI-E HMI Applications
To demonstrate real-world applications, the designed tactile sensor is utilized as an HMI interface to control virtual characters in the Google Dino Game and Super Mario Bros Game. Pressing different areas triggers corresponding actions, such as jumping and crouching in Google Dino, while varying press durations enable distinct motion amplitudes, like low- and high-altitude jumps in Super Mario Bros. Additionally, the tactile sensor controls a UR5e robotic arm, allowing it to grab objects at specified locations and arrange them sequentially. These demonstrations highlight the sensor’s ability to capture and transform subtle tactile inputs into real-time, intuitive controls for virtual environments and robotic systems. Supplementary Video 2 showcases real-time interactions with the Google Dino Game, Super Mario Bros Game, and the Universal Robots UR5e.
Figure 12: HMI applications (see supplementary video 2 for real-time control demonstrations).

### v2

@@ III-A3 Implementation
The voltage encoder $\mathcal{E}_{\text{vol}}$ and decoder $\mathcal{D}_{\text{vol}}$ are tailored with customized scales and hyperparameters for the $64\times 64$ E2IM input. The embedding dimension is set to 256 to effectively capture spatial information. An asymmetric architecture enhances feature extraction by using 12 Transformer layers in $\mathcal{E}_{\text{vol}}$ and a simpler decoder with 2 layers in $\mathcal{D}_{\text{vol}}$. All layers employ 4 attention heads to ensure consistency.
@@ III-A4 E2IM’s Enhancement on Pretraining
The MAE study [32] showed that mask sampling strategies strongly affect model performance, though the reasons were not deeply explored. On high-semantic ImageNet images [36], masking less semantic structural information made reconstruction easier.
In contrast, as shown in Fig. 1, the EIM is a low-semantic voltage representation. Large patch masking obscures key structural features, degrading reconstruction quality, while very small patches (e.g., $1\times 1$) permit easy interpolation but limit meaningful feature learning. Thus, for a $16\times 16$ EIM, the minimum optimal patch size $P$ is 2.
To address this, we upsampled EIM from $16\times 16$ to $64\times 64$ with $4\times 4$ patches in E2IM, improving feature learning capacity. Ablation studies (Table I) on 25,000 samples show E2IM reduces MSE by 99.95% and increases SSIM by 66.07% compared to EIM, confirming its superior representation quality.
@@ III-B1 Framework
$\arg\min_{\mathcal{E}_{tact},\mathcal{D}_{tact}}\mathcal{L}_{tact}=\mathbb{E}_{\textbf{t}}\left[\|\hat{\textbf{t}}-\textbf{t}\|^{2}\right],$ (10)
@@ III-B2 Implementation
The tactile autoencoding model $\mathcal{D}_{tact}(\mathcal{E}_{tact}(\textbf{t}))$, following VQ-VAE’s frame [38], is stacked with convolutional ResNet blocks. Each convolutional ResNet block consists of two ResNet layers followed by a convolution layer. These blocks perform downsampling or upsampling, with the number of channels in each layer ranging from [32, 64, 128].
@@
III-C Fine-tuning with limited voltage-tactile pairs
A key challenge in learning-based tactile sensing is the limited availability of annotated real-world data, due to the high cost and manual effort required to collect and annotate precise tactile measurements under controlled conditions. To overcome this, we fine-tune a unified model composed of a pretrained voltage encoder $\mathcal{E}_{\text{vol}}$ and tactile decoder $\mathcal{D}_{\text{tact}}$, using a small set of voltage–tactile pairs. This enables effective adaptation to the target task despite data scarcity.
Since we have small fine-tuning (i.e., down to 2,500 paired) and large pretraining datasets (i.e., 450,000 unlabeled), it often leads to over-fitting. To solve this problem, we propose strategies for few-shot fine-tuning for tiny datasets as follows:
@@ III-C1 Latent space compression
As outlined in Sections III-A and III-B, the deep neural network architecture of the autoencoder achieves a high compression ratio by encoding voltage and tactile data into a low-dimensional latent space. This compressed representation facilitates efficient data transformation and adaptation during the fine-tuning process. We keep 75% random masking in the MAE encoder for lower computation and real-time gains. Masking can be optionally removed for accuracy boosts (see Suppl. Mat.).
@@ III-C2 Partial tuning
Considering the significant differences in parameters between the voltage encoder $9.54M$ and the tactile decoder $1.23M$, as shown in Fig. 2, we freeze all parameters in the voltage encoder $\mathcal{E}_{vol}$ and only update the parameters in the tactile decoder $\mathcal{D}_{tact}$ during the fine-tuning stage, as shown in Fig.2. The latent variables from voltage measurement conditioning $\textbf{z}_{v}$ and tactile $\textbf{z}_{t}$ are connected via direct linear projection without any hidden layers:
@@
$\arg\min_{\mathcal{D}_{tact}}\mathcal{L}_{tact}=\mathbb{E}_{\textbf{z}_{t},\textbf{t}}\left[\|\mathcal{D}_{tact}(\textbf{z}_{t})-\textbf{t}\|^{2}\right],$ (15)
@@
Optionally, as illustrated in Fig. 2, Eq. (15) can be further improved by multiplying the sensor’s sensitivity matrix $S$:
$\arg\min_{\mathcal{D}_{tact}}\mathcal{L}_{tact^{+}}$ $=\mathbb{E}_{\mathbf{z}_{t},\mathbf{t},S}\Big[(1-\lambda)\|\mathcal{D}_{tact}(\mathbf{z}_{t})-\mathbf{t}\|^{2}$
$\quad+\lambda\|S\odot(\mathcal{D}_{tact}(\mathbf{z}_{t})-\mathbf{t})\|^{2}\Big]$ (16)
where $\lambda$ denotes a control weight, $\odot$ represents the element-wise multiplication (see Suppl. Mat.).
@@ IV-A Dataset
As shown in Fig. 3, the sensing region, measuring $100\,mm\times 100\,mm$, is divided into $48\times 48$ pixels with a background conductivity of 0.00312 S/m. Circular areas with diameters ranging from 7.5 mm to 27.5 mm are used to represent tactile interactions, emphasizing both position and intensity. The conductivity within these touch regions is randomly varied between 0.05 and 2 times the background value to simulate diverse touch effects. The dataset generated comprises 500,000 samples, with each of the five subsets containing 100,000 samples. Each sample includes one E2IM and corresponding Ground Truth (GT). These subsets correspond to different numbers of touch regions, ranging from 1 to 5 touch areas per sample. Each subset is further partitioned into training, validation, and testing datasets using an 18:1:1 ratio. This results in a total of 450,000 samples for training, 25,000 for validation, and 25,000 for testing. All data were generated using COMSOL Multiphysics and Matlab. The detail of the framework of the 2D EIT simulation dataset generation is shown in Supplementary Fig. 1. As shown in Supplementary Fig. 2, the waveform shapes of the reference voltage signals from simulation and real-world measurements are highly consistent.
@@ IV-C Quantitative evaluation
Quantitative Metrics RE PSNR $C$ MSE ($\times 10^{-2}$)
@@ V-A Sensor Fabrication
We designed a dual-conductivity layer tactile sensor combining a hydrogel layer and a CBG conductive layer, leveraging the complementary properties of each material. Hydrogels are widely utilized in electronic skin applications due to their desirable electrical properties, flexibility and biocompatibility [25]. The CBG layer with lower resistance improves EIT-based sensor performance. Fig. 6 shows the fabrication process and design of the EIT-based tactile sensor. The sensor dimensions are $120\,mm\times 120\,mm\times 10\,mm$. A step-by-step description of the fabrication process is available in the Supplementary Material.
@@
The dual-layer design combines the flexibility and biocompatibility of the hydrogel with the enhanced conductivity and structural integrity of the CBG layer, maximizing signal response under applied pressure and making it well-suited for applications in electronic skin and human-machine interfaces. To enhance durability and maintain long-term functionality, the entire hydrogel layer was encapsulated with a stretchable silicone elastomer to serve as a physical barrier against moisture loss [40].
@@ V-B Sensor Characterization
To evaluate the sensitivity of the developed sensor, we conducted an experiment where the sensor was subject to a continuously increasing localized pressure with strains ranging from 0 mm to 5 mm, as illustrated in Fig. 7. The results demonstrate a clear increase in the conductivity distribution as the compressive strain increases. Conductivity distribution images were generated for compressive strains of 1 mm, 2 mm, and 3 mm. At a compressive strain of 1 mm, the changes in conductivity are subtle yet detectable, while at 2 mm, the effect becomes more pronounced. These results demonstrate the sensor’s sensitivity and capability for accurate and responsive tactile detection.
@@
To evaluate the sensor’s durability and repeatability, we performed a cycling test involving 300 pressing cycles with compressive strains from 2 mm to 5 mm. The results show that the sensor maintains stable output throughout the cycles, with consistent trends in signal changes (see Supplementary Fig. 3). During this test, a dynamic baseline approach was adopted to mitigate slight signal drifts attributed to the lagging effect of the sensor after prolonged pressing. These results confirm the sensor’s robustness and suitability for applications requiring long-term, repeated tactile interactions.
All experiments in this study were conducted using the three-axis linear robot (IGUS, DLE-RG-0001), which ensured precise control over the position and depth of the presses (see Supplementary Fig. 4). All experiments were carried out under ambient laboratory conditions, without additional environmental control, and the pressing was performed at a constant strain rate of approximately 1 mm/s. An in-house developed EIT system was utilized for data acquisition as described in [33].
@@ VI-A Real-world Dataset
A diverse real-world dataset was created using 84 custom-fabricated touch tips produced by a 3D printer. The design of the touch points, including their positions and sizes, was randomly generated, with pressing depths precisely controlled between 2 mm and 5 mm.
We collected a total of 5700 annotated samples, distributed as follows: 613 samples with 1 touch point, 1,063 with 2 touch points, 1,043 with 3 touch points, 1,004 with 4 touch points, 995 with 5 touch points, 493 with annular touch and 489 with L-shaped touch. Each sample consists of a 104-dimensional vector of voltage measurements, which is transformed into a $64\times 64$ E2IM matrix and a $48\times 48$ image representing the press map. The samples were randomly shuffled and split into three datasets in a 7:2:1 ratio for training, validation, and testing. In addition, we augmented the training set with 30 frames of pure noise (i.e., no-touch) data to better simulate sensor noise and improve model robustness. To expedite training, we reused the pre-trained models from the simulation phase for both the PTET algorithm and the baseline DNN algorithm. This approach is justified because the pretraining model only captures the fundamental features of voltage measurements and pressure maps.
@@ VI-B Quantification of Reconstruction
Quantitative Metrics RE PSNR $C$ MSE ($\times 10^{-2}$)
@@ VI-C Performance on Complex Touches
To evaluate the model’s effectiveness in tactile reconstruction during practical operations, we tested it with a series of complex touch scenarios. These included configurations involving multiple fingers, annular touch, annular touch with finger, and L-shaped touches.
Fig. 9 presents a comparison among three methods: the $l1$ regularization reconstruction[42], SOTA DNN, and the proposed PTET framework. The results highlight the following observations: In the multi-finger touch phantom(Fig. 10a), both $l1$ and the DNN failed to resolve clear touch boundaries, while PTET reconstructed distinct, spatially localized regions corresponding to each finger. For the annular touch cases(Fig. 9b and 9c), the $l1$ and the DNN results showed partial or broken ring structures. In contrast, PTET successfully captured the complete annular shape, including the subtle structural difference introduced by the central finger in Fig. 10c. In the L-shaped touch phantom (Fig. 9d), $l1$ reconstruction was severely blurred, and the DNN yielded only partially correct contours. However, PTET produced a sharply defined L-shape, demonstrating superior spatial resolution and generalization to previously unseen topologies.
These results indicate that the PTET framework’s ability to generalize effectively to diverse real-world touch interactions even without specific labels. These results demonstrate the model’s adaptability and robustness for practical applications. Supplementary Video 1 shows the real-time tactile reconstruction performance on selected phantoms.
@@ VI-D Comparison of Different Training Datasets
$C$ 0.8633 0.4349 0.4455
MSE ($\times 10^{-2}$) 0.1551 0.9069 0.8829
@@
VI-E Human Machine Interaction (HMI) Applications
As shown in Fig. 11, to demonstrate real-world applications, the designed tactile sensor is utilized as an HMI interface to control virtual characters in the Google Dino Game and Super Mario Bros Game. Pressing different areas triggers corresponding actions, such as jumping and crouching in Google Dino, while varying press durations enable distinct motion amplitudes, like low- and high-altitude jumps in Super Mario Bros Game. Additionally, the tactile sensor controls a UR5e robotic arm, allowing it to grab objects at specified locations and arrange them sequentially. These demonstrations highlight the sensor’s ability to capture and transform subtle tactile inputs into real-time, intuitive controls for virtual environments and robotic systems. Supplementary Video 2 showcases real-time interactions with the Google Dino Game, Super Mario Bros Game, and the Universal Robots UR5e.
Figure 11: HMI applications (see Supplementary Video 2 for real-time control demonstrations). (a) Google Dino Game. (b) Super Mario Bros Game. (c) UR5e robotic arm.